# GraphLoRA: Empowering LLMs Fine-Tuning via Graph Collaboration of MoE

### v1

@@ Abstract
Low-Rank Adaptation (LoRA) is a parameter-efficient fine-tuning method that has been widely adopted in various downstream applications of LLMs. Together with the Mixture-of-Expert (MoE) technique, fine-tuning approaches have shown remarkable improvements in model capability. However, the coordination of multiple experts in existing studies solely relies on the weights assigned by the simple router function. Lack of communication and collaboration among experts exacerbate the instability of LLMs due to the imbalance load problem of MoE. To address this issue, we propose a novel MoE graph-based LLM fine-tuning framework GraphLoRA, in which a graph router function is designed to capture the collaboration signals among experts by graph neural networks (GNNs). GraphLoRA enables all experts to understand input knowledge and share information from neighbor experts by aggregating operations. Besides, to enhance each expert’s capability and their collaborations, we design two novel coordination strategies: the Poisson distribution-based distinction strategy and the Normal distribution-based load balance strategy. Extensive experiments on four real-world datasets demonstrate the effectiveness of our GraphLoRA in parameter-efficient fine-tuning of LLMs, showing the benefits of facilitating collaborations of multiple experts in the graph router of GraphLoRA.
@@ 1 Introduction
Recently, increasing attention has been paid to the Parameter-Efficient Fine-Tuning (PEFT) strategies of Large Language Models (LLMs). For example, the Low-Rank Adaptation (LoRA) technique with LLMs has been widely adopted to downstream applications of LLMs due to the lower computational costs (Hu et al. 2021). Although the efficiency problem has been addressed in PEFT methods (e.g., LoRA) by updating fewer parameters, their performance inevitably declines, especially in multi-task scenarios. According to the scaling law, researchers attempt to utilize Mixture-of-Expert (MoE) to enhance the overall capability of LLMs (Shazeer et al. 2016). The research that applies MoE in PEFT of LLMs attracts increasing attention due to the remarkable improvements of model performance in practical downstream applications (Gou et al. 2023; Wang and Wu 2024; Yang et al. 2024b; Li et al. 2024a).
One practical way to incorporate MoE in PEFT LLMs is using LoRA aligned with the Feed Forward Network (FFN) layer in a transformer block (Li et al. 2024a; Wang et al. 2022; Dou et al. 2023a). The weight of each expert is assigned by a simple softmax function (termed as router function or gating function) with the information from input tokens. In these cases, an expert in the MoE framework is unable to explicitly share information with other experts, which limits their collaboration capability and exacerbates the instability of LLMs due to the imbalance load problem of MoE. The model quickly converges to a localized optimization, leading to the limited utilization of the comprehensive potential capabilities of MoE. Recent studies (Shazeer et al. 2016; Zoph et al. 2022; Luo et al. 2024; Dou et al. 2023a; Wang et al. 2022; Shazeer et al. 2016) add a balance loss that constraints the allocated frequency of each expert to alleviate the imbalance load problem of MoE. However, we believe that a more effective collaboration strategy in the router allocation process is emergency required to improve the coordination of multiple experts.
To address this issue, we propose a novel graph router for PEFT of MoE in LLMs, termed GraphLoRA. The graph router is conducted on a MoE Graph, which contains an input token node and all expert nodes. The information of the input tokens and collaboration signals from other experts is captured by graph neural networks (GNNs), enabling all experts to understand the input knowledge and share information with other experts effectively due to the shared parameters in GNNs. The graph router assigns weight to each expert according to the projection vector that is computed based on the representations of all expert nodes learned in GNNs. Then, Top-K experts with the highest assigned weights are activated to perform efficient sparse tuning in downstream applications. To further empower the capability of each expert and enhance their collaboration, we propose two novel coordination strategies: the Poisson distribution-based distinction strategy and the Normal distribution-based load balance strategy. Specifically, a Poisson distribution loss function is used to release the different capabilities of each expert, and a normal distribution is adopted to keep the balance of the activated frequency of each expert in a natural pattern.
Equipped with the effective collaboration mechanism of experts in the graph router and two coordination strategies, our GraphLoRA achieves the SOTA model performance in downstream applications. Our contributions are summarized as follows:
@@
  • We propose a novel GNNs-based MoE PEFT framework GraphLoRA to address the instability problem caused by the imbalance load of MoE in LLMs fine-tuning. A graph router assigns weights to experts with consideration of their collaborative interactions on the MoE Graph.
@@
  • We propose two novel coordination strategies, i.e., the Poisson distribution-based distinction strategy and the Normal distribution-based load balance strategy, to enhance the expert’s capability and collaboration among them.
@@
  • Extensive experiments conducted on four real-world datasets with three typical LLMs demonstrate the effectiveness of our GraphLoRA in terms of model Accuracy, Stability, and Efficiency, showing the benefits of empowering collaborations of multiple experts in PEFT of LLMs.
@@ 2.1 Mixture-of-Expert (MoE)
Empowered by the collaboration of multiple experts, the Mixture-of-Expert (MoE) framework has achieved remarkable performance in many applications (Chen, Jie, and Ma 2024; Li et al. 2024a; Lin et al. 2024; Li et al. 2024b; Zadouri et al. 2023), such as recommendation, video understanding, natural language generation, and so on. Each expert in the MoE framework specializes in handling a subset of input information. Different experts coordinate together to gain benefits for the overall MoE framework. According to the number of experts activated by the router function, the MoE methods can be classified into two categories, i.e., dense MoE and sparse MoE. In dense MoE, all experts are activated in the learning process, which enhances the model capability but suffers from high computational costs. In sparse MoE, only a selected subset of experts are activated, which not only reduces the computational overhead but also leverages their specialized knowledge for optimal results. The participation of each expert in the computation process is assigned by a router function (or gating function), ensuring an optimal blend of their specialized contributions. Specifically, in the MoE literature of LLMs (Zoph et al. 2022), the router is usually a Softmax function that controls the engagement of expert computations. All experts are allocated to computation in the dense MoE depends on the weights assigned by the router. In sparse MoE, only Top-K experts with the highest weights are selectively activated in the learning process. The MoE framework is widely used in the pre-training and fine-tuning process of LLMs. The pertaining of LLMs requires massive datasets and a high demand for computational resources. Our work uses open-sourced pre-trained LLMs, like Llama and Qwen, and focuses on utilizing the MoE to enhance the Parameter-Efficient Fine-Tuning (PEFT) of LLMs in specific domains.
2.2 Parameter-Efficient Fine-Tuning (PEFT)
@@
The aim of fine-tuning LLMs is to optimize the model’s performance in specific downstream tasks. Due to the high computational costs of updating all parameters in the full fine-tuning approach, the Parameter-Efficient Fine-Tuning (PEFT) technique is proposed to address this problem (Hu et al. 2021; Houlsby et al. 2019; Li and Liang 2021; Lester, Al-Rfou, and Constant 2021). In PEFT, only a small subset of parameters are updated of the base LLMs, making it more efficient and practical in real applications. The PEFT methods can be generally classified into three types, i.e., LoRA (Hu et al. 2021), adapter tuning (Houlsby et al. 2019), and prompt tuning (Lester, Al-Rfou, and Constant 2021). Adapter tuning uses feed-forward up and feed-forward down projection matrices in the transformer block and only updates two adapter metrics in the fine-tuning process. Prompt tuning introduces a set of learnable prompt embeddings that are appended to the input tokens. The most widely used PEFT technique is LoRA, which uses a low-rank decomposition technique and updates the decomposed parameter matrix in the model training.
@@
Recent studies improve the model performance by integrating the MoE framework with PEFT in LLMs fine-tuning literature (Wu, Huang, and Wei 2024; Zadouri et al. 2023). In the transformer block of LLMs, each expert is a LoRA module working on either the FFN layer (Dou et al. 2023b; Wang et al. 2022; Li et al. 2024a), attention layer (Luo et al. 2024; Gou et al. 2023; Zhu et al. 2023), the whole transformer block (Zadouri et al. 2023; Liu et al. 2023; Gao et al. 2024) and each layer (Wu, Huang, and Wei 2024). Then the router function of MoE is designed to blend the contributions of all experts. Among them, the adoption of LoRA in the FFN layer attracted the most attention due to its extensive applicability in various LLM tasks in recent years.
@@
Our work aims to explore the effective MoE collaboration mechanism in the parameter-efficient fine-tuning of LLMs area. The MoE consists of multiple LoRA components that work on the FFN layer in the transformer block. Different from existing MoE studies, in which all experts coordinate by the traditional router function (Softmax function), we design a graph router in our GraphLoRA framework, which leverages graph neural networks to learn the collaboration information among all experts in the MoE graph.
@@
Figure 1: The overview of MoE architectures in the FFN layer. (a) shows the LoRA component applied in the FFN layer of the transformer block. (b) is the typical MoE architecture with LoRA in LLMs with a router (Softmax) function to assign weights to each expert. (c) is our proposed GraphLoRA architecture with a graph router based on the MoE graph. For different input information, distinctive experts are optimized by the Poisson distinction loss function. For all input information, the activated frequency of each expert is balanced by the normal distribution loss.
@@ 3 Preliminary
3.1 Low-Rank Adaption (LoRA)
Low-rank adaption (LoRA) (Hu et al. 2021) is a widely used parameter-efficient fine-tuning (PEFT) technique in pre-trained large language models (LLMs). LoRA adopts a low-rank decomposition technique to update the parameters of the decomposed parameter matrix to learn the data distribution of the specific downstream task. It works on the feed-forward layer (FFN layer) in a transformer block of LLMs. For a linear layer in FFN represented as $\mathbf{h}=\mathbf{W}\mathbf{x}$, the update of the decomposed parameter matrix of LoRA is defined as:
@@ 3.1 Low-Rank Adaption (LoRA)
where $\mathbf{x}\in\mathbb{R}^{I}$ is the representation of input information, and $\mathbf{W}\in\mathbb{R}^{{O}\times{I}}$ is the pre-trained parameter matrix of LLMs. $\mathbf{A}\in\mathbb{R}^{r\times I}$ and $\mathbf{B}\in\mathbb{R}^{O\times r}$ are the low-rank matrix in LoRA with $r\ll min(I,O)$. $\alpha$ represents the magnitude of the changes in $\mathbf{W}$. Only decomposed matrix $\mathbf{A}$ and $\mathbf{B}$ are updated in the fine-tuning process.
@@ 3.2 Mixture-of-Expert (MoE-LoRA)
where $\mathcal{E}_{i}$ is the $i$th expert network and $N$ is the number of experts. The router function $\mathcal{R}(\cdot)$ is a Softmax function that generates a probability distribution for the weights assigned to all experts.
4 GraphLoRA
The overview architecture of GraphLoRA is shown in Fig. 1 (d). This section introduces the details of the graph router in GraphLoRA and the propagation process in the FFN layer. Besides, to empower all experts to fully utilize their unique capabilities and keep the load balance, we propose two novel coordination strategies: load balance strategy and expert distinction strategy to enhance the GraphLoRA capability.
@@ 4.1 Graph Router
In typical MoE-LoRA literature, all experts coordinate solely depending on the router function, which is a simple Softmax function that assigns the weights for each expert. The inefficient collaboration of all experts exacerbates the imbalance load problem of MoE in LLM fine-tuning. To enhance the collaboration of all experts, we propose a graph router to replace the simple router function. The graph router function in GraphLoRA assigns weights to each expert in collaboration with other experts on the MoE graph.
Specifically, given the MoE graph $G=(V,E)$. The node set $V=\{e_{1},e_{2},...e_{N},x\}$ consists of all expert nodes and the input-token information node $x$ (after the self-attention layer and normalized layer in traditional transformer block). The edges between nodes are randomly constructed and controlled by an edge density hype parameter $\beta$. For the input-token node $x$ and its neighborhood expert nodes $N_{e}(x)$ in the MoE graph, we use a graph neural network (GNN) to learn their interaction information, which not only implies the learning capability of each expert to the input-token node but also captures the collaboration information among all experts. The feed-forward process in GraphLoRA can be formulated as:
@@
where $\textit{GNN}(\cdot)$ is the graph neural network that learns the representation of expert node $e_{i}$ using information from its neighbors $N(e_{i})$ that contains other experts and input token node $x$. $\mathcal{F}(\cdot)$ is the project function that maps the representations from all experts into one unique vector. $\mathcal{R}(\cdot)$ is the Softmax function that assigns probability weights to all experts.
Following the sparse-gated strategy in MoE studies (Zoph et al. 2022; Li et al. 2024a), which had been proposed to address the computational overhead problem in LLMs. We use the Top-K experts in the feed-forward process of the FFN layer, defined as:
@@
The graph router in GraphLoRA takes full advantage of the collaborative information aggregations by GNNs on the MoE graph. This alleviates the localized convergence tendencies caused by a small number of experts activated in the existing MoE studies (Shazeer et al. 2016). Instead of proving a collaboration mechanism for all experts, we propose two novel coordination strategies, i.e., the Poisson distribution-based expert distinction strategy and the normal distribution-based load balance strategy in GraphLoRA, to enhance the model capability.

### v3

# GMoE: Empowering LLMs Fine-Tuning via MoE Graph Collaboration

@@ Abstract
The sparse Mixture-of-Experts (MoE) architecture of large language models (LLMs) confronts an inherent issue of load imbalance arising from the simplistic linear router strategy, which ultimately causes the instability and inefficient learning of LLMs. To address this challenge, we introduce a novel MoE graph-based framework GMoE, aimed at enhancing the collaboration among multiple experts. In GMoE, a graph router function is designed to capture the collaboration signals among experts. This enables all experts to dynamically allocate information derived from input data by sharing information with their neighboring experts. Moreover, we put forward two coordination strategies in GMoE: the Poisson distribution-based distinction strategy and the Normal distribution-based balance strategy, to further release the capacity of each expert and increase the model stability in the fine-tuning of LLMs. Specifically, we leverage a parameter-efficient fine-tuning technique, i.e., Low-Rank Adaptation (LoRA), to implement the graph MoE architecture. Extensive experiments on four real-world benchmark datasets demonstrate the effectiveness of GMoE, showing the benefits of facilitating collaborations of multiple experts in LLM fine-tuning. The code of experimental implementation is available at https://github.com/BAI-LAB/GMoE
GMoE: Empowering LLMs Fine-Tuning via MoE Graph Collaboration
Ting Bai Beijing University of Posts and Telecommunication baiting@bupt.edu.cn                        Yue Yu Beijing University of Posts and Telecommunication loadingyy@bupt.edu.cn                        Le Huang Beijing University of Posts and Telecommunication lehuang@bupt.edu.cn
Zenan Xu Tencent AI Lab zenanxu@tencent.com                        Zhe Zhao Tencent AI Lab nlpzhaozhe@tencent.com                        Chuan Shi Beijing University of Posts and Telecommunication shichuan@bupt.edu.cn
@@ 1 Introduction
The Mixture-of-Expert (MoE) architecture has emerged as a promising approach to enhance the overall performance of large language models (LLMs) under the scaling law theory. The application of MoE in the fine-tuning of LLMs has drawn significant attention due to the substantial improvements it brings to model performance in practical downstream applications. In a typical MoE setup, a simple linear function, acting as the “router function” or “gating function,” assigns weights to each expert based on the information from input tokens, then the top-k experts with largest assigned weights are activated in the fine-tuning process. However, sparse MoE methods face significant challenges that impede their performance and introduce model instability, with a core issue being their linear router strategy that creates severe load imbalance, where a few experts are over-trained while others are under-utilized Fedus et al. (2022).
Recent studies have introduced load balance losses to constrain expert allocation frequencies, thereby alleviating MoE load imbalance (Fedus et al., 2022; Shazeer et al., 2016; Zoph et al., 2022; Luo et al., 2024; Dou et al., 2023a; Wang et al., 2022). However, these approaches primarily focus on regularization-based constraints for load balancing while failing to address a critical limitation: the absence of effective communication and collaborative mechanisms among experts during router allocation. Without expert collaboration, the router relies solely on individual input patterns instead of leveraging collective capabilities. This coordination gap exacerbates inefficiencies: even with balanced allocations, the lack of inter-expert communication prevents dynamic optimization to exploit each expert’s unique strengths. As a result, experts remain underutilized, limiting the model’s potential and leading to instability during LLM fine-tuning.
These challenges highlight a critical research gap: while load regularization alleviates imbalance, enhancing the stability and efficiency of LLM fine-tuning with MoE requires moving beyond standalone constraints to design architectures that enable effective collaboration among experts. To address this issue, we propose GMoE, a novel graph-based MoE framework that explicitly models collaboration among experts. Specifically, GMoE introduces a graph router to encode collaborative signals among experts via a MoE Graph integrating input tokens and expert nodes. Using graph neural networks (GNNs), the router aggregates information iteratively, capturing both token features and inter-expert interactions. This enables experts to jointly process inputs and dynamically share knowledge, leading to more informed, collaborative routing decisions.
To further empower the capability of each expert and enhance their collaboration, we propose a novel coordination strategy: the Poisson distribution-based distinction strategy. Specifically, the Poisson distribution loss function promotes specialization by encouraging experts to handle distinct input aspects, thereby stimulating their unique capabilities. Concurrently, the Normal distribution-based balance strategy regulates activation frequencies, naturally balancing the overall workload distribution. This synergistic combination enables experts to discriminatively process inputs while maintaining workload balance, fostering coordinated interaction and enhancing overall model efficiency.
We adopt the parameter-efficient fine-tuning (PEFT) technique, i,e., Low-Rank Adaptation (LoRA), to enable an efficient implementation of our graph MoE architecture for LLM fine-tuning. Through the expert collaboration mechanism in the graph router and two coordination strategies, our GMoE framework achieves state-of-the-art performance and remarkable stability during the fine-tuning of diverse benchmark datasets. Our contributions are summarized as follows:
@@
  • We propose a novel GNNs-based MoE framework GMoE to address the instability problem of MoE in LLMs fine-tuning. A graph router assigns weights to experts with consideration of their collaborative interactions on MoE Graph.
@@
  • We propose a novel coordination strategies, i.e., the Poisson distribution-based distinction strategy. Along with the Normal distribution-based load balance strategy, our graph based MoE architecture, GMoE, fully unleash the capabilities of individual experts and significantly enhance their collaboration.
@@
  • Extensive experiments conducted on four real-world benchmark datasets with three typical base LLMs demonstrate the effectiveness of our GMoE in terms of model Accuracy, and Stability, showing the benefits of empowering collaborations of multiple experts of LLMs.
@@ 2.1 Mixture-of-Expert (MoE)
Empowered by the collaboration of multiple experts, the Mixture-of-Expert (MoE) framework has achieved remarkable performance in many applications (Chen et al., 2024; Li et al., 2024a; Lin et al., 2024; Li et al., 2024c; Zadouri et al., 2023). Each expert in the MoE framework specializes in handling a subset of input information and different experts coordinate together to gain benefits for the overall MoE framework. According to the number of experts activated by the router function, the MoE methods can be classified into two categories, i.e., dense MoE and sparse MoE (Dou et al., 2023b; Li et al., 2024b). In dense MoE, all experts are activated in the learning process, which enhances the model capability but suffers from high computational costs. In sparse MoE, only a selected subset of experts are activated to use their specialized knowledge and achieve optimal results, thereby reducing computational overhead. The participation of each expert in the computation process is assigned by a router function (or gating function), ensuring an optimal blend of their specialized contributions (Zoph et al., 2022; Tang et al., 2025). Specifically, the router is usually a simple linear function that controls the engagement of expert computations. In the dense MoE, all experts are allocated to computation depending on the weights assigned by the router. In sparse MoE, only Top-K experts with the highest weights are selectively activated in the learning process (Shazeer et al., 2016).
For example, the typical sparse MoE approach Switch Transformers (Fedus et al., 2022) routes only a single expert for the input token to reduce the computation costs, but it faces the load imbalance problem due to lacking collaboration of multiple experts. Existing studies have incorporated auxiliary losses (Bengio et al., 2015; Fedus et al., 2022) to maintain a balanced load among experts in sparse MoE architectures. However, exploring an optimal collaboration mechanism among the activated experts to enhance their load distribution remains an ongoing challenge in this field. In our work, we propose a novel graph router to assign weights to experts with consideration of their collaborative interactions on the MoE Graph, so as to achieve better capability and stability in fine-tuning of LLMs.
2.2 PEFT with MoE
@@
The aim of fine-tuning LLMs is to optimize the model’s performance in specific downstream tasks. Due to the high computational costs of updating all parameters in the full fine-tuning approach, the Parameter-Efficient Fine-Tuning (PEFT) technique is proposed to address this problem (Hu et al., 2021; Houlsby et al., 2019; Li and Liang, 2021; Lester et al., 2021; Tian et al., 2024). In PEFT, only a small subset of parameters are updated of the base LLMs, making it more efficient and practical in real applications. The most widely used PEFT technique is LoRA (Hu et al., 2021), which uses a low-rank decomposition technique and updates the decomposed parameter matrix in the model training.
@@
Recent studies have shown improvements of model performance by integrating the MoE framework with PEFT in LLMs fine-tuning literature (Wu et al., 2024; Zadouri et al., 2023). In the transformer block of LLMs, each expert is a LoRA module working on either the FFN layer (Dou et al., 2023b; Wang et al., 2022; Li et al., 2024a), attention layer (Luo et al., 2024; Gou et al., 2023; Zhu et al., 2023), the whole transformer block (Zadouri et al., 2023; Liu et al., 2023; Gao et al., 2024) and each layer (Wu et al., 2024). Then a router function of MoE is designed to blend the contributions of all experts. Among them, the adoption of LoRA in the FFN layer attracted the most attention due to its extensive applicability in various LLM tasks in recent years.
@@
Our work focuses on exploring the effective MoE collaboration mechanism to enhance the PEFT of LLMs in downstream tasks. The MoE consists of multiple LoRA components that work on the FFN layer in the transformer block. Different from the conventional router function employed in existing MoE studies, where the weight assigned to each expert is solely dependent on its own input information, we design a novel graph router, which leverages graph neural networks to learn the collaboration information among all experts based on a MoE graph.
@@
Figure 1: The overview of MoE architectures in the FFN layer. (a) The LoRA component is applied in the FFN layer of the transformer block. (b) The typical MoE architecture with LoRA in LLMs with a linear router function to assign weights. (c) Our proposed GMoE architecture with a graph router based on the MoE graph. For different input information (input1 and input2), the distinctive capability of experts is optimized by the Poisson distinction loss. For all input information, the activated frequency of each expert is balanced by the normal distribution loss.
@@ 3 Preliminary
3.1 Low-Rank Adaptation (LoRA)
Low-rank adaptation (LoRA) (Hu et al., 2021) is a widely used parameter-efficient fine-tuning (PEFT) technique in pre-trained LLMs. LoRA adopts a low-rank decomposition technique to update the parameters of the decomposed parameter matrix to learn the data distribution of the specific downstream task. It works on the feed-forward layer (FFN layer) in a transformer block of LLMs. For a linear layer in FFN represented as $\mathbf{h}=\mathbf{W}\mathbf{x}$, the update of the decomposed parameter matrix of LoRA is defined as:
@@ 3.1 Low-Rank Adaptation (LoRA)
where $\mathbf{x}\in\mathbb{R}^{I}$ is the representation of input information, and $\mathbf{W}\in\mathbb{R}^{{O}\times{I}}$ is the pre-trained parameter matrix of LLMs. $\mathbf{A}\in\mathbb{R}^{r\times I}$ and $\mathbf{B}\in\mathbb{R}^{O\times r}$ are the low-rank matrix in LoRA with $r\ll min(I,O)$. $\alpha$ represents the magnitude of the changes in $\mathbf{W}$. Only decomposed matrices $\mathbf{A}$ and $\mathbf{B}$ are updated in the fine-tuning process.
@@ 3.2 Mixture-of-Expert (MoE-LoRA)
where $\mathcal{E}_{i}$ is the $i$th expert network and $N$ is the number of experts. The router function $\mathcal{R}(\cdot)$ employs a linear function to produce a probability distribution from input information, which serves as the weights allocated to all experts. Following the sparse-gated strategy in MoE studies (Zoph et al., 2022; Li et al., 2024a), which had been proposed to address the computational overhead problem in LLMs. The top-K experts with the highest weights are selectively activated.
4 GMoE
The overview architecture of GMoE is shown in Fig. 1 (d). This section introduces the details of the graph router in GMoE and the propagation process in the FFN layer. Besides, to empower all experts to fully utilize their unique capabilities and keep the load balance, we use two coordination strategies: load balance strategy and expert distinction strategy to enhance the GMoE capability.
@@ 4.1 Graph Router
In typical MoE-LoRA literature, all experts coordinate solely depending on the simple linear router function. The weight assigned to each expert solely depends on the input information, with no explicit collaboration or communication mechanisms among the experts. This may result in the over-training of only a few experts and under-training of others stemming from its simplistic router strategy. The inefficient collaboration of all experts exacerbates the imbalance load problem of MoE in LLM fine-tuning.
To enhance the collaboration of all experts, we propose a graph router to replace the simple router function. The graph router function in GMoE assigns weights to each expert in collaboration with other experts on the MoE graph. Specifically, given the MoE graph $G=(V,E)$. The node set $V=\{e_{1},e_{2},...e_{N},x\}$ consists of all expert nodes and the input-token information node $x$ (after the self-attention layer and normalized layer in traditional transformer block). The input node is connected to all the expert nodes to ensure all experts can receive the input information equally. The edges between all expert nodes are randomly constructed and controlled by an edge density hype parameter $\beta$. For the input-token node $x$ and its neighborhood expert nodes $N_{e}(x)$ in the MoE graph, we initialize their features using the commonly adopted Glorot uniform initialization. Subsequently, we use a graph neural network (GNN) to learn their interaction information, which not only implies the learning capability of each expert to the input-token node but also captures the collaboration information among all experts.
The feed-forward process in GMoE can be formulated as:
@@
where $\textit{GNN}(\cdot)$ is a two-layer graph neural network that learns the representation of expert node $e_{i}$ using information from its neighbors $N(e_{i})$ that contains other experts and input token node $x$. $\mathcal{F}(\cdot)$ is a projection function, and $\mathcal{R}(\cdot)$ is the Softmax function that assigns probability weights to all experts.
Then we use the Top-K experts in the feed-forward process of the FFN layer, defined as:
@@
The router in GMoE takes full advantage of the collaborative information aggregations by GNNs from other expert nodes. This alleviates the load unbalance and localized convergence tendencies caused by a small number of experts activated in the existing MoE studies (Shazeer et al., 2016). Except for introducing a collaboration mechanism of all experts, we use two coordination strategies, i.e., the Poisson distribution-based expert distinction strategy and the normal distribution-based load balance strategy in GMoE, to further release the capacity of each expert and increase the model stability.